OpenStudy (anonymous):

find the integral of (2x+)/(x-1). please help me with this question

OpenStudy (anonymous):

\[\int\frac{2x+\cdots}{x-1}~dx\]

OpenStudy (anonymous):

I'm assuming you're just missing a constant, in which case you can make a substitution. Let \(u=x-1\), so \(x=u+1\), and so that \(du=dx\): \[\int\frac{2(u+1)+\cdots}{u}~du\]